### v2

Abstract
This paper explores the potential for leveraging Large Language Models (LLM) in the realm of online advertising systems. We delve into essential requirements including privacy, latency, reliability as well as the satisfaction of users and advertisers which such a system must fulfill. We further introduce a general framework for LLM advertisement, consisting of modification, bidding, prediction, and auction modules. Different design considerations for each module is presented, with an in-depth examination of their practicality and the technical challenges inherent to their implementation. Finally, we explore the prospect of LLM-based dynamic creative optimization as a means to significantly enhance the appeal of advertisements to users and discuss its additional challenges.
1 Introduction
In the vast landscape of online search engines, the role of advertisements has become pivotal, shaping the digital experience for users globally. The enormity of the market, with trillions of dollars at play, underscores the economic significance of advertising. In a world where people increasingly expect free access to information and services, advertisements play a crucial role in subsidizing these offerings. The democratization of education, information, and other resources is made possible through ad-driven models, enabling individuals to not only consume content for free but also contribute to the digital ecosystem. Additionally, the symbiotic relationship between advertisements and content creation creates a feedback loop, fostering economic growth. Several online platforms often offer a subscription model for people to disable the advertisement in using their services, however, the advertisement market is becoming increasingly dominant, e.g., Netflix recently started a ad-supported plan Netflix (2022).
Recently, large language models (LLMs) have gained widespread adoption among users, serving various functions including question answering, content generation, translation, code completion, and more Brown et al. (2020); Nijkamp et al. (2022); Fried et al. (2022); Wang and Komatsuzaki (2021); Liu et al. (2023). The proliferation of AI-driven assistant language models, such as ChatGPT, has contributed to a growing trend wherein individuals increasingly use such models to address their inquiries, occasionally replacing traditional search engines as their primary information-seeking tool. The substantial usage volumes stemming from diverse users shave prompted companies offering these tools, which we call LLM providers, to contemplate revenue generation through advertising AdWeek (2023); Crunch (2023); Microsoft (2023). Consequently, an interesting and fundamental question arises:
How can LLM providers generate revenue through running an online advertisement?
The concept of online advertising has been extensively studied within the realm of search engines, where auctions are conducted among advertisements from advertisers when a user inputs a query. We focus on the prospect of transposing this online advertising model and auction framework to the context of large language models, exploring the potential challenges in this context. In this paper, our primary focus revolves around discussing different criteria for LLM advertisement and evaluating different frameworks based on them.
Figure 1: An example of providing unstructured advertisement in the LLM. Left column refers to the case where we ask above question from ChatGPT 3.5, and to incorporate the ads we use queries specified in the right column.
We start with illustrative scenarios where a user asks a technical question about computers (Figure 1). Without advertisement, LLM would typically generate a response to address the user’s query. On the other hand, within the content of the generated response, there exists an opportunity for the LLM provider to strategically incorporate advertisements for computer manufacturers. There is a spectrum of possibilities for including ad content, ranging from ads alongside the output, explicitly putting advertisements in the output like display ads, and including advertisements within the generated output. Also, in case the advertisements are included within the generated output, there are several possibilities such that the ads may simply replace one of the element in case they are written as a list of elements (e.g., Figure 4, denoted by structured output), or the ads itself is incorporated within the original text/figure without hurting/changing the context of the original output (e.g., Figure 1, denoted by unstructured output).
We will focus on the case where the ads are incorporated in the unstructured output of the LLM, since the other cases could be handled using techniques from modern online search/display advertisement system. Moreover, this kind of advertisement can be more interesting for both users and advertisers due to the huge capabilities of language models in generating high-quality responses while promoting products and services from advertisers. [^2] Indeed, it’s worth noting that Microsoft Bing has been running such types of advertisements in its AI chatbot Microsoft (2023). Then, a natural question is how the system can incorporate advertisements into the generated output to satisfy both the advertisers and the user. Further, it is not trivial how the advertisers should communicate with the system to bid on each modified output.
In this paper, we discuss several desiderata for operating an online advertisement over LLM and propose a general framework to meet the requirements. Our framework consists of several modules: (i) modification in which the original output of LLM is modified; (ii) bidding that advertisers utilize to bid on the modified outputs; (iii) prediction in which the LLM provider computes required information about advertisements; and (iv) auction in which the advertises compete and the final output is selected. We discuss different designs and technical challenges of each of those modules, along with other technical challenges to implement the system in a practical manner.  [^3] We first focus on presenting a single advertisement in the LLM output, however, generalization of the proposed framework to incorporate multiple ads at once is discussed in Appendix B.2. Finally, in Section 6, as a LLM’s (more generally generative model’s) usecase of independent interest, we discuss how LLM can empower the online advertisement system to tailor the content of ads itself to individual users, leveraging their specific contexts to generate more appealing content, which is called as dynamic creative optimization (DCO) or responsive ad in the literature.
2 Related works
Here we present related works in terms of LLM, online advertisement system, and their intersection. Further related works can be found in Appendix A.
Large Language Models
Advancements in AI, NLP, and conversational agents, driven by Transformer architecture (Vaswani et al., 2017), have given rise to models like GPT-3 (Brown et al., 2020) and BERT (Devlin et al., 2019). These models revolutionize chatbots, enabling context-aware, human-like interactions across diverse domains Abd-Alrazaq et al. (2020); Nicolescu and Tudorache (2022).
Online Advertisement
Online advertising, particularly within the context of sponsored search auctions, has evolved in recent years, with notable contributions from prior research. Sponsored search auctions have been a subject of extensive investigation, emphasizing the optimization of bidding strategies and keyword relevance. [^4] Online advertisement can be categorized by either of search or display ads. We here focus on search ads. Edelman et al. (2007) provided valuable insights into the economics of sponsored search auctions, shedding light on the intricate dynamics of keyword auctions. Goel et al. (2009) proposed a contract auction between the advertiser and the publisher, and introduce impression-plus-click pricing for sponsored search auction as an application. We refer to the book by Roughgarden (2010) for more detailed discussion.
Mechanism design with LLM agents
Duetting et al. (2023) recently proposes a mechanism design problem when the bidders are LLM agents. More specifically, the bidders bid on desired distributions of the next token, given the history of the tokens. They denote this problem by token auction, and provide theoretical foundations of what a reasonable mechanism should look like. Indeed, their model can be considered as a version of our advertiser modification and dynamic bidding framework. Notably, they require the advertisers to provide their desired distribution and corresponding bidding for every token, which may require significant communication burden to run the mechanism in practice. In contrast, the advertisers in our framework only bid once for the modified output, thereby inducing less communication cost. [^5] In Section 5.4 and Appendix B, we provide further discussion on Duetting et al. (2023) to compare it with our framework and how one can practically adapt their mechanism design perspective in online advertisement system within our framework.
3 Requirements
In this section, we outline our expectations for an LLM advertisement system, which should be evaluated based on the following criteria:
Privacy
In any advertisement system where users and advertisers interact, it is imperative to maintain the confidentiality of users’ privacy. In essence, all aspects pertaining to a user’s context, information, and data must be kept secure, as any inadvertent disclosure can pose significant risks.
Reliability
LLM advertisement systems need to associate with advertisers, making the entire system vulnerable to harmful behaviors from them. To mitigate that, LLM providers should consider all aspects of the behavior of the advertisers, ensuring the reliability and alignment of the system (Hendrycks et al., 2020).
Latency
Users expect rapid interaction with LLM service, e.g., ChatGPT typically generates prompt responses, without any noticeable delay. Incorporating advertisement in LLM output, inevitably, will add some latency to the overall system. Nevertheless, this added delay should be minimal, ensuring that it does not significantly disrupt the user experience.
User satisfaction
When an advertisement is added to LLM output, it’s crucial to maintain a high level of quality in the generated content. Users do not wish to encounter an excessive amount of advertising in the generated responses, especially when it doesn’t align with their search or interests. Advertisement involvement should not significantly degrade LLM output, as doing so would lead to reduced user satisfaction, thereby decreasing the user retention. For instance, in modern online advertisement system, this is often handled using floor prices to filter out irrelevant ads to maintain the quality of the user experience.
Advertisers satisfaction
As advertisers pay LLM providers to involve advertisements in their outputs, they expect appropriate exposure of ads in them. Whether their products, services, or any other form of advertising, they expect these items to be showcased in a compelling and engaging manner within the output. This presentation should be interesting for users, ultimately driving revenue for the advertisers.
LLM provider’s revenue
Obtaining revenue is the primary goal of providing LLM advertisement services, thus, LLM providers should ensure that the total additional cost of advertisement is compensated by the revenue obtained from advertisers. It is worth noting that advertisements may potentially reduce the overall number of users engaging with the system, which could have adverse effects on LLM providers.
4 Framework for LLM Advertisement
Figure 2: Overall framework of LAS.
In this section, we present overall framework for LLM advertisement system (henceforth LAS). We introduce several feasible types of architectures for LAS, each of which has its own advantages and disadvantages that will be discussed together. Mainly, we focus on a scenario in which a single user provides a single query $q$ to the LLM, and the original output by LLM is given by $X$. Further, a context $c$ captures a variety of features that are relevant to the advertisement recommendation, e.g., history of the previous queries, user segment, region, and date. Although the number of advertisers (bidders) varies from time to time, at the moment the user inputs the query $q$, we suppose that there are $n$ advertisers indexed by $V=\{\text{adv}_{1},\text{adv}_{2},\ldots,\text{adv}_{n}\}$. For simplicity, we assume that each advertiser $\text{adv}_{i}$ is equipped with a single advertisement (ad) he/she wants to post. We denote $\text{adv}_{i}$’s ad by $\text{ad}_{i}$ for $i\in[n]$, and let $D=\{\text{ad}_{1},\ldots,\text{ad}_{n}\}$.
Overall, we divide the LAS as follows based on their functionalities [^6] These functionalities may be distributed to multiple market players such as supply side platform, display side platform, or ad exchange, as does in the current online ad eco-system. into $4$ modules: (i) output modification, (ii) bidding, (iii) prediction, and (iv) auction. This overall framework is illustrated in Figure 2.
In short, the user asks a query $q$ to the system along with a context $c$, and the LAS first generates the original output $X$. Then, the modification module generates specific output $X_{i}$ for each of the advertisers, and the bidding module generates corresponding bid $b_{i}$. At the same time, the prediction module predicts the quality of the output $X_{i}$ by measuring the user satisfaction rate $\text{sr}_{i}$ and click-through-rate $\text{ctr}_{i}$, which will be elaborated later. Finally, the auction module runs auction to determine the final output and corresponding payment amount to charge the advertiser. Also, note that we do not restrict the final output to be exactly belonging to one of $X_{i}$, but can be totally different one based on the advertisers’ bids $b_{i}$ and corresponding output $X_{i}$. For the rest of the section, we will explain each module and its responsibility/functionality in a sequential manner.
Modification module
Modification module takes responsibility for generating modified output based on corresponding ads. This module takes the pair of $(q,X,c)$ and a set of advertisements $D$ as the input, and returns $(X_{i})_{i\in[n]}$ where $X_{i}$ denotes the modified output for $\text{ad}_{i}$.
Overall, we consider two approaches based on who generates the modified outputs. 777 For either cases, one can generate the modified output by throwing an additional query to the LLM as presented in Figure 1.
  1. 1.
    In the advertiser modification model, the role of generating the modified output is delegated to each advertiser, and collect the modified outputs from them.
  2. 2.
    In the LAS modification model, LAS directly generates the modified output.
Bidding module
Bidding module generates bid based on the modified outputs. The input to the bidding module is the query $q$, context $c$, and modified outputs $(X_{i})_{i\in[n]}$, and the output will be a bid $(b_{i})_{i\in[n]}$, where $b_{i}$ is the bid for $ad_{i}$. Similar to the generation model, we consider two models for the bidding module.
  1. 1.
    In the dynamic bidding model, we deliver relevant information, e.g., query $q$, context $c$, the original output $X$, [^8] To protect the user’s privacy, one may consider delivering an encrypted context $\hat{c}$ or the one that excludes private information. along with the modified output $X_{i}$ to the bidder for each query, and then the bidder returns the corresponding bid.
  2. 2.
    In the static bidding model, each bid is determined by some keywords based on some pre-committed contract, without any additional communication with the advertiser.
Since the static bidding model may operated solely based on some keywords that are relevant to the advertisement products, we often refer this to as keyword bidding. In Section 5, we discuss several extensions of these bidding models, thereby allowing more flexibility for advertisers to customize their advertisement strategy by bidding.
Dynamic bidding model may be of interest to the advertisers who have their own measures to estimate the quality of the modified output and thus want to customize their bid in a way that might not be captured by the static bidding. For example, in extreme, suppose that LAS may only consider click-through-rate of the modified output and does not care about the user satisfaction rate or how much the inserted advertisement relates to the context of the output. In this case, the advertiser may want to customize their bid by measuring the quality of the output by itself, since he may be afraid of the system inserting its ads in an awkward manner.
Prediction Module
Prediction module is responsible for computing user’s satisfaction-rate (SR) and click-through-rate (CTR). User’s SR indicates how much the user is satisfied with the given output. This SR affects the final decision making process of LAS since if there is a chance that the user is very disappointed with the modified output, then it should reconsider displaying that output as it may hurt the user experience a lot. On the other hand, CTR indicates the probability that the user clicks the ad link included within the modified output. This is also critical in determining the auction winner since this directly affects the revenue of the LAS. More precisely, if an advertiser’s bidding method is cost-per-click (CPC), then CTR proportionally affects the advertisement revenue as the expected revenue of transferring its ad can be computed as CPC times the CTR. Overall, both of SR and CTR can be defined as a function of original output $X$, modified output $X^{\prime}$, query $q$, and context $c$, which returns a real value in $[0,1]$. Detailed methodologies for estimating/learning these functions will be discussed in Section 5.
Auction module
Computed all the required parameters above, we finally run the auction module to determine who wins the auction and how much the advertiser will be charged. The input to the auction module is the tuples of $(\text{bid}_{i},\text{sr}_{i},\text{ctr}_{i})_{i\in[n]}$ which denotes bid amount, satisfaction rate, and click-through-rate, respectively. Output of the auction module is an allocation $a\in\{0,1\}^{n}$ and a payment $p\in\mathbb{R}_{\geq 0}$. Precisely, auction module determines the auction format which consists of the allocation function and the payment function, where the allocation function decides which advertisement to deliver and the payment function computes how much the advertisers pay LAS. [^9] In its current form, this does not allow the possibility of adjusting the final output in order to balance between multiple advertiser’s preferred output. Our framework, however, does not preclude possibility for such extension, which we will elaborate more in Appendix B.
Importantly, the main goal of LAS is to maximize its revenue in a long term manner. To this end, this should balance a trade-off between short-term revenue and the user retention. Therefore, the objective can be modeled as a function from bid amount, CTR, and SR to a nonnegative real number that indicates a score of a modified output, i.e., selecting $i^{*}=\operatorname*{argmax}_{i\in[n]}\text{Obj}(\text{sr}_{i},\text{ctr}_{i},\text{bid}_{i})$. We do not discuss how to choose the objective function here, given the vast literature on the sponsored search auction. After designing the score function, one should determine an auction format. Also, Given the long line of literature on auction design, one can pick the most preferred mechanism, e.g., first-price auction (FPA) or second-price auction (SPA).
5 Challenges
Recall that our overall framework consists of four modules: modification, bidding, prediction, and auction module. For each module, we discuss several characteristics and technical challenges in implementing the proposed architectures in practice, evaluating their overall practicality with respect to the criteria defined in Section 3. Further discussions can be found in Appendix B due to page limit.
5.1 Modification module
Comparing modification models
In the advertiser modification model, LAS should somehow deliver $q,X,C$ to each $\text{adv}_{i}$. This may leak some private information of the user to the bidders, thereby suffering privacy issues, which is critical in terms of the user experience. Partially disclosing the information may help resolve the issue.
In addition, the advertiser modification model has reliability issue since there is a possibility that the advertiser-modified output includes some illegal or spamming contents. The LAS may devise an additional module to maintain its robustness against such kind of adversarial behavior, however, this is still a cost of the LAS to bring them in the system. This further possibly degrades the user satisfaction since the advertiser-generated output may not be satisfactory for the user compared to what the LLM generates. Furthermore, this requires additional communication between the LAS and the advertisers, thereby increasing the overall latency. Overall, advertiser modification model may not be feasible in practice due to several issues discussed above.
On the other hand, LAS modification model does not suffer the privacy, reliability and latency issues. Furthermore, this model would be more in favor of the user since LAS would basically hope to improve the user experience. However, one disadvantage is that this may not fully reflect the advertiser’s preference on the output modification process, thereby partly decreasing the advertiser’s satisfaction. This issue can be resolved by enabling more flexibility in the architecture, which will be discussed in what follows.
Beyond advertiser and LAS modification
In order to improve the advertiser’s satisfaction for the LAS modification module, one approach is to allow the advertiser to submit indirect indicators that represent its preference on what it expects the modified output to be. More in a detail, once the user asks a query $q$, LAS delivers $q$, the original output $X$, and possibly the context $c$ to the advertiser, and the advertiser returns a string $Y$ that indicates its preference in words. After then, LAS may apply $Y$ as a prompt in the LLM to generate the modified output. This approach allows more flexibility for the advertiser to express its preference on the output. Note that, however, this still requires additional communication cost compared to the LAS modification, which possibly increases latency.
5.2 Bidding module
Comparing bidding models
The biggest advantage of the dynamic bidding model is that, it would induce higher advertiser’s satisfaction rate, thereby attracting more advertisers to participate. This is because the advertisers have full flexibility in deciding their bids after observing the modified output. Similar to the advertiser modification model, dynamic bidding model may have privacy issue if LAS discloses some private information to the advertiser. Moreover, dynamic bidding model may suffer additional latency issue since the entire set of modified output should be delivered to the advertisers. Note that there is no reliability concern as only the bid amount is communicated.
Static bidding model, however, does not suffer any of the privacy, latency, reliability issues. However, as the advertisers cannot fully customize their bid per each query based on the modified output, this may degrade the advertiser satisfaction.
Beyond static and dynamic bidding
In the static bidding model, a natural approach to improve the advertiser’s satisfaction is to enable a more generic static function as the contract between the advertiser and LLM provider. For example, in a simplistic approach, the advertiser may specify specific user segments to deliver their advertisements, e.g., children under 16, or may screen out some traffics based on geographical information, e.g., users far from national parks. More critically, since the advertiser may want to provide a tailored bid with respect to the quality of its modified output, LAS can propose the bidding contract as a function between the quality of the modified output to a non-negative real number. For instance, the bidding function on the committed contract can be a linear function that is proportional to the distance measure between the modified output and the original output, e.g., charge \$$\$5\cdot d$ if the similarity distance is $d$. Note that in this case, the advertiser may need to partly understand how LAS estimates the distance measure and how it’s defined. More generally, LAS might construct more refined way to measure the units of interest, attention, relevance for the modified output in terms of $\text{adv}_{i}$, and let the advertisers commit to either of the contract based on its taste.
5.3 Prediction module
The process of estimating CTR will be analogous to that in the modern online advertisement system. We can train a prediction system to predict $\text{ctr}_{i}\in[0,1]$ given the input $X_{i},q,c$ for each $i\in[n]$ based on the feedback data from the user. [^10] In practice, one may extract some useful features from each $q,X$ and $c$, and learn the mapping from features to the true value. Feedback data will be the direct action of the user indicating whether the user clicks the ad or not. This prediction system can be trained in an online manner to update the estimate in (almost) real time, which will increase the accuracy of the prediction.
To estimate SR, one may first consider an explicit approach given an assumption that the original output is optimal. If we assume that the original output is optimal, then essentially a distance between the original output $X$ and the modified output $X_{i}$ will be a good measure to predict the SR since if $X_{i}=X$, then the user will be the most satisfied. In this context, given a string $s(X)$ that represents the output $X$, we propose that the marginal probability ${\mathrm{Pr}\left[\,s(X)|q\,\right]}$ as an indicator for SR. Precisely, we define an output distance by $d(X,X_{i}):=d({\mathrm{Pr}\left[\,s(X)|q\,\right]},{\mathrm{Pr}\left[\,s(X_{i})\,\right]}|q)$ where $d$ denotes some valid metric. Note that the probability for an output to be obtained can be computed via standard method in the literature (e.g., see Vaswani et al. (2017)). Further, one can consider more general distance function $d$ to improve the accuracy of the prediction, such as functions to measure the semantic similarity of two phrases (Mikolov et al., 2013; Cer et al., 2018; Conneau et al., 2019).
Alternatively, we can learn SR via feedback data from the user. A useful indicator of whether the user is satisfied with the output is whether the user re-generates the output given the same (or similar) query. Hence, one may aim to learn a function which outputs the sr, given the query $q$, modified output $X^{\prime}$, and context $c$.  [^11] For instance, as is typical in the modern online advertisement system, one may use factorization machine based online algorithm after extracting some features from user, context, and the queries. More detailed discussion can be found in McMahan et al. (2013). This approach does not assume that the original output $X$ is indeed optimal, thereby allowing the possibility that the user may be satisfied with a modified output $X_{i}$ even though its distance from $X$ is measured to be far. However, one should implement an independent training system to operate this learning process, leading to additional burden for LAS. To incorporate advantages of these two approaches, one can consider an online learning model where the distance is also integrated as one of the features to predict it, i.e., $d(X,X_{i})$ is used as a feature of the SR prediction system. In case $d(X,X_{i})$ indeed has some positive correlations with the true user satisfaction rate, this would increase the accuracy of the prediction.
5.4 Auction module
Multiple ads at once
Recall that in our auction module, the modified output is fully determined based on the allocation function so that if it is allocated to $\text{adv}_{i}$, then the modified output $X_{i}$ will be displayed to the user. One approach to generalize this framework is, to let the final output to be displayed the one that does not necessarily belong to $\{X_{i}\}_{i\in[n]}$, but rather a new output $X^{\prime}$ that somehow interpolates $\{X_{i}\}_{i\in[n]}$. [^12] We discuss further comparison to the token auction model by Duetting et al. (2023) in Appendix B.1. Let us denote the new output $X^{\prime}$ by balanced output. By doing so, it might be possible to deliver multiple advertisements in a fair manner, [^13] Another way to display multiple ads at once will is discussed in Appendix B.2 under our original framework. thereby allowing LAS to bring more revenue by charging multiple advertisers at once. One subtle issue is that, since each advertiser bids $b_{i}$ on delivering $X_{i}$, they may not want to bid the same amount for the balanced output $X^{\prime}$, thereby degrading the advertiser’s experience on participating in the LAS. In the static bidding model, as discussed in Section 5.2, this can be handled by committing to a contract based on measures that represents the advertiser’s preferences more in a refined manner. In the dynamic bidding model, of course, this may simply be handled by incorporating additional steps of querying the bid for the final output again.
6 Dynamic Creative Optimization via LLM
Figure 3: Integrating personalized advertisement in the images generated by ChatGPT 4. In the left panel, a user requests car recommendations within the price range of \$40,000 to \$60,000, and the chat bot, when capable of advertising, generates a “BMW 3 Series” image (left). Personalized ads modify this image for a 60-year-old woman (middle) or man (right) based on user context like age and gender. In the right panel, a user seeks machine learning conferences in Florida. The chat bot advertises “SunShine Travel” with its logo on a generic Florida image (left). With personalized ads, the model tailors promotions, offering flight tickets from the user’s location and showing more relevant Florida images. For example, for a young unmarried person from DC, the output advertises DC to Florida flights (middle) and features an image of beach there, while for a married person in California, it promotes California to Florida flights with images of amusement parks. This enhances user experience with targeted and appealing content.
Now, we discuss about the possibility of advanced dynamic creative optimization by using LLM. Dynamic creative optimization (DCO)  [^14] This is often called as dynamic product ads (DPA), but DPA typically refers to the process of selecting which item to advertise from a given catalog of an advertiser, whereas DCO denotes the process of modifying the ad itself., in the traditional search and display ad markets, refers to the technique of dynamically adjusting the combination of ad assets and constructing an advertisement that best fits the customer’s preference. It can markedly enhance the quality and relevance of the ad to the user by tailoring the contents of ad in a personalized manner. Many online advertisement platforms adopt DCO-based ads, which is often called as responsive ad GoogleDisplayAd ; GoogleSearchAd .  [^15] As traditional DCO is restricted to dynamically construct the complete image file of ad, we here restrict our attention to the image-based ads as well. However, it is apparent that LLM based DCO would work beyond image-based ads.
In the canonical DCO framework, a static ad consists of a single image file that can only be displayed in a specific set of ad channels, whereas dynamically optimized ad consists of a single ad template and multiple options for each asset that fills in the ad template. This is typically done by the advertiser registering multiple options for each asset given an ad template, and then the system picks the optimal combination. This procedure aims at increasing the CTR and SR, which would eventually improve the welfare efficiency of the online ad ecosystem McMahan et al. (2013) as well as the revenue of the ad platform.
Unlike from the traditional DCO techniques, the arise of LLM, especially those support the output with image such as ChatGPT 4, enables more involved process of dynamic creation of ads, as well as replaces the role of traditional DCO in online ad platforms. For example, given a user query and context, the platform may send a query to LLM to customize the image of ad to attract the user by capturing user preference. We provide representative examples in Figure 3 where two scenarios of personalizing ads is depicted.
Incorporating responsive ad
To integrate responsive ads, a more advanced modification module is imperative. This module should not only enable the alteration of the original response to include ads but also incorporate them in a manner that takes into account user preferences. This entails utilizing user contexts, e.g., gender/location/the current device used for the query, in the ad generation process. To be more specific, the modification module should incorporate user context in the process of modifying the original response. This makes ads more appealing to the user by considering their preferences. Leveraging language models, we can prompt them to factor in various elements related to user context during the generation of modified outputs. As illustrated in Figure 3, information about users’ context can be utilized to create more captivating advertisements for them. Further, the prediction module may interact with the modification module to commit to the modified output that mostly increases user’s experience, i.e., CTR and SR.
Cost sharing model
Unlike from the traditional DCO technique that simply determines the efficient combination of ad assets, LLM-based responsive ad can do much beyond by creating truly new content. On the other hand, such process necessarily entails increased use of computational resource, especially since more queries to LLM will be required to customize the ads. We here discuss several plausible cost sharing models that LAS might adopt.
The simplest one will be to charge the advertisers whenever the contents are dynamically modified. For example, the advertiser and LAS commit to a mutual contract on how much the advertiser pay the LAS for each responsive ad, and LAS charges the advertiser whenever it responsively modifies the ad contents. As the advertiser might hope to use responsive ad only if the user context is sufficiently relevant to its ads, the contract might specify when the advertiser’s ads will be responsively changed. Also, LAS may provide several options for advertiser to determine how frequently and largely it should responsively modify the ads.
This model would work well for those charged by cost-per-impression, however, advertisers with cost-per-click or cost-per-conversion may not willingly commit to the contract since they desire to pay the platform only if the actual click or conversion happen, even though the platform indeed responsively modified its ads for every impression. LAS might give several options to the advertisers under which event they will be charged to pay for responsive ad. Ads with cost-per-click model may still want to pay for responsive ads per impression or per click. For the latter, LAS may offer a larger amount of commission per click by predicting the average CTR of the corresponding ads.
LAS must also decide whether or not to account the increased amount of payment for each advertiser as the increment of the actual bid or just regard it as a separate commission. For the former case, the advertisers who are willing to charge more for the responsive ad commission fee will take advantage of getting higher chance of delivering their ads, compared to those who do not use responsive ad option. By doing so, the auction module can select the ad that truly maximize the total revenue expected from delivering a single advertisement, thereby increasing the platform’s revenue.
Finally, we remark that these involved process demand additional computational resources, potentially impacting latency when incorporating dynamic ads. Online advertisement typically requires the overall latency of real time bidding process to be within $100$ms, however, we experience several second of delay to modify the image output presented in Figure 3. Thus, it is technically interesting to balance the tradeoff between the ad’s quality and latency, e.g., naively by using cache, and also it might be possible to pass along part of such cost to the advertisers via monetary charge.
7 Conclusion
This paper addresses the opportunities and challenges for leveraging an online advertisement system with LLM. We first discuss essential requirements which a reasonable LLM advertisement system should satisfy. To fulfill such desiderata, we introduce several frameworks and substantiate their validity. We compare proposed frameworks with their advantages and limitations. Then we discuss further technical and practical challenges in designing efficient LLM-based online advertisement system. Finally, as a usecase of independent interest, we present the potential of advanced dynamic creative optimization by using LLM.
8 Acknowledgements
This work is partially supported by DARPA QuICC, NSF AF:Small #2218678, and NSF AF:Small #2114269, and a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Meta grant 23010098, HR00112090132 (DARPA/RED), HR001119S0026 (DARPA/GARD), Army Grant No. W911NF2120076, NIST 60NANB20D134, the NSF award CCF2212458, an Amazon Research Award and an award from Capital One. The authors thank Sébastien Lahaie for valuable discussions.
Appendix A Further Related Works
Chatbot advertisement
With emergence of chat-bot in various websites, several companies started to run a chat-bot advertisement, which is often referred to as sponsored message. In the chat-bot advertisement, the companies send some private messages on the chat-bot conversation to advertise some products. This is very different from operating an online advertisement system over LLM, since chat-bot advertisement simply sends some messages, but our model captures the scenario in which the user asks a query and then LLM provides a output including some advertisements. Notably, Microsoft Bing started to operate an online advertisement with Bing chat-bot (e.g., see articles  AdWeek (2023); Crunch (2023); Microsoft (2023)). Indeed, once the user inputs a query in the chat-bot, the output often includes advertisements with explicit flags that indicates this is an advertisement.
Dynamic creative optimization
Dynamic Creative Optimization (DCO) is an online advertising strategy that involves dynamically tailoring the content of an ad based on user behavior, demographics, and user context Baardman et al. (2021). Instead of displaying a static, one-size-fits-all ad, DCO allows advertisers to create and serve personalized and relevant ad content to different audience segments. By doing so the platform can improve the likelihood of user engagement, click throughs, and conversions. DCO typically involves the use of technology and algorithms to automate the process of selecting and assembling ad elements, such as images, copy, and calls-to-action, in real-time. This customization is done based on user data, collected either in real-time or from previous interactions. Unlike DCO has been typically used only in display advertising not search advertising Choi et al. (2020) since the tailored ad creatives, e.g., personalized hyperlinks to websites, would not have much impact for structured output of standard search engines. Due to the unstructured nature of LLM advertisement, the use of techniques similar to dynamic creative optimization would be highly likely to be present in future. Our paper suggests the use of LLM, more generally generative AI, to personalize the advertisement itself to improve the relevance and quality of the advertisement.
Figure 4: An example of providing structured advertisement in the LLM. Left column refers to the case where we ask above question from ChatGPT 3.5, and to incorporate the ads we use queries specified in the right column.
Appendix B Further Discussion
B.1 Comparison to Duetting et al. (2023)
The approach we discussed at the end of Section 5.4 is similar to the token auction model proposed in Duetting et al. (2023), under which the advertisers equipped with their own LLM directly provide a desired distribution over the subsequent token and corresponding bid, and then the mechanism finally decides the final distribution by running a token auction. They provide a methodology to aggregate distribution and commits to the final one, while constructing a reasonable payment scheme to respect the reasonable properties of auction in the mechanism design perspective. Importantly, however, their model does require oracle access to the LLM of the advertisers to generate each token, thereby requiring more communication cost to construct the entire output. Our advertiser modification and dynamic bidding model can be thought as a version of their model, but the advertiser only bids on the final deterministic output. In this context, one may consider generalizing this model by letting the advertisers to bid on a distribution over the final output, thereby allowing the advertisers to represent their preferences more in a refined way. Note, however, that since the entire space of possible output may be extremely large, let the advertisers bid a distribution over such space might be practically/computationally infeasible. Alternatively, one can consider a version in which each advertiser provides a set of outputs, (partial) ordering over it with corresponding bid, and then the auction module determines the final output based on designed mechanism.
B.2 Further challenges and practical concerns
Parallelizing the process and adapting multiple types of advertisers
Recall that our overall architecture consists of (i) modification, (ii) bidding, (iii) prediction, and (iv) auction modules. Although each of the process should be serialized in order to handle a single advertiser’s process, it is important to observe that the process (i)-(iii) is totally independent between the advertisers. Hence, one efficient to way to minimize the latency is to process steps (i) to (iii) in an asynchronous manner between the advertisers. To fit the desired latency requirement, one may consider excluding some advertisers such that the steps (i) to (iii) are not completed within some maximal delay. After collecting all the feasible parameters, LAS can pass the required parameters to step (iv) and run the auction.
One additional advantage of this procedure is that, one can easily adapt to multiple types of advertisers. In practice, several advertisers may want to dynamically bid or modify the output by themselves, while others with small businesses may not want to take too much of responsibilities. For instance, consider the following scenario.
  1. 1.
    $\text{adv}_{1}$ wants to modify the output by itself, and dynamically bids on it.
  2. 2.
    $\text{adv}_{2}$ does not want to modify the output by itself, but wants to dynamically bid.
  3. 3.
    $\text{adv}_{3}$ does not want to modify the output by itself, and wants to bid in a static manner.
If one conceptually separates the steps (i), (ii) and (iii) between the advertisers, we can adapt all of these advertisers by running a separate procedure for each advertiser, and then finally runs the step (iv).
Displaying multiple ads simultaneously
One common feature in the modern sponsored search auction is that, it usually display multiple ads upon a single query, which is mostly to maximize the search engine’s revenue. Given the architecture we proposed, it is not straightforward to allow this feature, so we here discuss several approaches to deal with it.
First, consider a simple case in which all advertisers use LAS modification and static bidding model. In this case, one can consider every combination of advertisements as a subset of $[n]$, and then generate modified output with corresponding bid. Further, if there is an advertiser with dynamic bidding, LAS can simply delegate the bid generating process to that advertiser for every combination that includes this advertiser’s ad.
Obviously, these approaches are not computationally feasible if one consider every possible combinations of the subset. Therefore, the technical challenge lies in how one can optimally decide a subcollection of the entire power set to balance the trade-off between computational burden with corresponding latency and revenue of the auction.